Jakathy is 32 years older than her nephew, Eric. Five years ago, her age was seventeen times as old as his. What is Jakathy's age?
Answer by mananth(16946)

eric's age be x
Jakathy's age = x+32
..
5 years ago their ages
Eric = x-5 years
Jakathy = x+32-5=x+27 years
..
x+27 = 5(x-5)
x+27=5x-25
4x=52
x= 13 years Eric's age
Jakathy's age = x+32= 13+32=45 years